# Unveiling Ruby: Insights from Stack Overflow and Developer Survey

### v1

@@ Abstract.
Ruby is a widely used open-source programming language, valued for its simplicity, especially in web development. Despite its popularity, with over one million users on GitHub, little is known about the issues faced by Ruby developers. This study aims to investigate the key topics, trends, and difficulties faced by Ruby developers by analyzing over $498,000$ Ruby-related questions on Stack Overflow (SO), followed by a survey of $154$ Ruby developers. We employed BERTopic modeling and manual analysis to develop a taxonomy of $35$ topics, grouped into six main categories. Our findings reveal that Web Application Development is the most commonly discussed category, while Ruby Gem Installation and Configuration Issues emerged as the most challenging topic. Analysis of trends on SO showed a steady decline. A survey of $154$ Ruby developers demonstrated that $31.6\%$ of the participants find the Core Ruby Concepts category particularly difficult, while Application Quality and Security is found to be difficult for over $40\%$ of experienced developers. Notably, a comparison between survey responses and SO metrics highlights a misalignment, suggesting that perceived difficulty and objective indicators from SO differ; emphasizing the need for improved metrics to better capture developer challenges. Our study provides insights about the challenges Ruby developers face and strong implications for researchers.
@@ 1. Introduction
This study offers valuable insights into the main challenges that Ruby developers face, paving the path for library contributors and tool developers to create resources that effectively address these pain points, ensuring improvements in the Ruby ecosystem that align with developer needs. We open-source our collected dataset and codes to enable the reproducibility of the results. [^1] https://osf.io/d9qan/?view_only=9f19ef0dafae46b4bef8c857862ec35d
@@ 2. Related Works
Novelty of our work. While the current literature covers developers’ challenges in other programming languages and provides insights into Ruby’s popularity, community engagement, and collaborative practices, it does not address the specific challenges faced by Ruby developers. Our study fills this gap by analyzing Ruby-related SO questions and conducting a targeted survey of Ruby developers to identify key areas of difficulty and evolving interests.
@@ 3. Methodology
$\alpha$ (Equation 1) measures how relevant is a specific tag in $T$ to Ruby and $\beta$ (Equation 2) measures how significance is a specific tag in $T$. Tags meeting or exceeding defined thresholds for both heuristics were considered significantly relevant to Ruby. We applied thresholds from previous studies (Yang et al., 2023), setting six thresholds for $\alpha$ = {$0.05$, $0.1$, $0.15$, $0.2$, $0.25$, $0.3$} and six for $\beta$ = {$0.005$, $0.01$, $0.015$, $0.02$, $0.025$, $0.03$}. The first author evaluated all $36$ configurations to determine which configuration contained the tags most relevant to Ruby development. The best configuration we selected is $\alpha$ = $0.1$ and $\beta$ = $0.01$, aligning with the findings of earlier research (Yang et al., 2023). This refinement yielded the following tag set, $T$ = {ruby, devise, ruby-on-rails-5, rspec, rubygems, sinatra, activerecord, ruby-on-rails-3, nokogiri, rvm, ruby-on-rails, ruby-on-rails-4, hash}.
@@
Step 3: Data preprocessing. We preprocessed our dataset in multiple stages. First, we removed posts that their title and body had less than $50$ characters ($4,398$ posts were removed) For the remaining posts, we concatenated titles and bodies since our manual analysis showed both contained essential keywords. To create embeddings we removed punctuation marks and stop words from NLTK list (Bird et al., 2009). For creating the top $10$ words for each topic, we did another preprocessing which consisted of removing escape characters, HTML tags, links, code blocks, numbers, stop words, and punctuation marks, and lower-casing all the characters. These steps ensure we have a meaningful list of words for each topic and are used in previous studies (Haque et al., 2020; Yang et al., 2023; Bagherzadeh and Khatchadourian, 2019; Ahmed and Bagherzadeh, 2018; Zhao et al., 2024).
@@
To construct the taxonomy, we used a combination of GPT-4o and a human-in-the-loop approach, a methodology used in prior studies (Shah et al., 2023). The process began by providing GPT-4o with the titles and bodies of $15$ randomly selected questions from each of the $35$ topics, as well as the topic names. GPT-4o was instructed to generate a one-level taxonomy, aiming to create $15$ to $20$ middle categories based on this information and assign names to each of them. Next, two authors manually reviewed the titles and bodies of the $525$ questions (i.e., $15$ per topic), as well as the topic names. After closely examining the taxonomy generated by GPT-4o, the authors refined the groupings, ultimately producing $18$ middle categories. In the final step, we repeated this process to develop the highest level of the taxonomy. GPT-4o was asked to generate $5$ to $10$ higher-level groupings based on $18$ middle categories and $15$ randomly selected questions from each of them (i.e., $270$ posts), which would serve as the main categories. As a result, six main categories were created. By the end of the process, all Ruby-related questions were organized into one of these six main categories, which the authors manually evaluated.
@@
Step 6: Topic analysis. To assess the popularity and difficulty of topics, we adopted methods from prior research (Yang et al., 2023; Haque et al., 2020). Popularity is evaluated using two metrics: the average number of views for questions within each topic and the average score of questions in each topic. The average number of views shows how many times that question was viewed by people on SO. When a question is useful to the community, users can upvote that question and otherwise, downvote it. The score of each question is calculated by summing the number of upvotes minus the number of downvotes. So, if a question has a high average score it means that the question was very useful to the community. Following previous studies (Yang et al., 2023), our main metric for popularity is the average number of views. The difficulty is assessed through three distinct metrics that were used in previous studies (Yang et al., 2023; Haque et al., 2020). First, we considered the percentage of questions within each topic that do not have an accepted answer. Second, we evaluated the percentage of questions that remain unanswered. Third, we measured the median time required for questions in each topic to receive an accepted answer. Following previous studies (Yang et al., 2023), our main metric for understanding difficulty is the percentage of questions without an accepted answer. The author of the question can mark one of the answers as accepted answer. Therefore, if a high percentage of questions in a topic do not have accepted an answer, we can understand that the topic is difficult. Additionally, we analyzed the trends related to Ruby questions on SO over the years to understand how they have evolved.
@@
This study was approved by the Behavioural Research Ethics Board of the University of British Columbia.
@@
To validate the mechanics of our survey instrument, we conducted a pilot study with seven graduate students at ANONYMOUS University. These participants had experience with Python but not Ruby. Given the similarities between the two languages, we designed the pilot survey to mirror the final survey format while focusing on Python instead of Ruby (Chandran and Abraham, 2020). This approach allowed us to assess the survey’s clarity, structure, and usability without depleting the pool of potential Ruby respondents for the main study. While we recognize that Python and Ruby developers may face distinct challenges, the primary goal of the pilot study was not to compare the two languages but to ensure the survey was well-structured and functional. The pilot also helped identify potential issues and estimate the completion time. By testing the survey with Python developers, we established a solid foundation for the main study, which would later target Ruby developers. After that, we distributed the full survey via Prolific [^4] https://www.prolific.co, a platform previously used in similar research studies (Reid et al., 2022; Russo, 2022). To ensure data quality, participants were pre-screened for proficiency in Ruby and prior programming experience. Since the survey took approximately $10$ minutes to complete, an attention-check question was included, requiring a correct response for the participant’s submission to be considered valid. The survey was published on Prolific in August 2024, yielding a total of $211$ responses, of which $154$ were approved. Approved participants were automatically paid at a rate of £6 per hour following Prolific’s guidelines. To analyze the survey responses, as (Liang et al., 2024), we report the percentage of participants who selected each option for every topic.
@@ 4.1. RQ1: Topics and Taxonomy
Only $2.59\%$ of the questions were related to Software Architecture and Performance. The questions in this category are related to enhancing the efficiency and design of Ruby applications, design pattern queries, and improving modularity across multiple Rails applications.
@@ 4.2. RQ2: Analyzing the Topics
Average for all Ruby questions 2092.46 2.49 41.6 9.68 48 Total: 494321 -
@@
To understand whether the more difficult topics are also more popular, we conducted a Kendall rank correlation test at a $95$% confidence level (Hussain and Mahmud, 2019), adopted from prior research (Ahmed and Bagherzadeh, 2018). The null hypothesis ($H_{0}$) is that there is no significant relationship between average popularity and average difficulty and the alternative hypothesis ($H_{1}$) is that there is a significant relationship between average popularity and average difficulty. The Kendall correlation coefficient was found to be $-0.33$, indicating a moderate negative correlation between popularity and difficulty. However, the $p$-value of $0.47$ suggests that this correlation is not statistically significant. Thus, we cannot confidently assert that a meaningful relationship exists between average popularity and average difficulty.
@@
Figure 3. The trends of six main categories over time. The blue line represents the total number of category-specific questions asked over time, while the yellow line illustrates the percentage of category-specific questions relative to all Ruby-related questions.
@@ 4.4. RQ4: Survey Results-Time
On the right side of the chart, we can see the percentage of respondents who reported spending more than $1$ hour to resolve their problems, while the left-hand side reflects the percentages of those who managed to resolve their issues in less than $30$ minutes. The findings indicate that within the Software Architecture and Performance category, $29.5\%$ of participants required more than one hour to address their issues, which is the highest percentage of any category for those taking more than $2$ hours at $13.4\%$. In contrast, the Application Quality and Security category revealed that $44.9\%$ of participants resolved their issues under $30$ minutes. Furthermore, the Web Application Development category exhibited the highest percentage of participants, at $16.7\%$, who were able to resolve their problems in less than $15$ minutes.
@@
This discrepancy between RQ3 and RQ4 results highlights that developers perceive difficulty and time to resolution differently from what SO data indicates. This could be due to the data from SO and the survey being inherently different and not comparable. Another indication could be similar to the well-known “stated vs. revealed preferences” in behavioral sciences (De Corte et al., 2021); suggesting a potential need for alternative or additional metrics to better capture the complexity and difficulty of issues from a developer’s perspective and from SO. As our model shows, the variation in difficulty level is not well explained by time-to-resolution, meaning that time-to-resolution alone may not accurately reflect the difficulty that developers experience. Given the low $R^{2}$ value in our model, this finding likely points to a need for improved or supplementary metrics on SO, as current metrics seem insufficient for capturing perceived difficulty levels.
@@ 5. Discussions
The difference we found in RQ3 and RQ4 results of SO vs survey data, can be attributed to several key factors that highlight the inherent differences in how both data sources measure difficulty. Survey data is subjective, relying on developers’ individual perceptions that are influenced by personal experiences and prior knowledge among others. In contrast, SO data is based on objective, measurable metrics that capture the broader activity on the platform. As a result, comparing the two becomes challenging, as they measure fundamentally different aspects of the developers’ experiences.

### v2

@@ Abstract.
Ruby is a widely used open-source programming language, valued for its simplicity, especially in web development. Despite its popularity, with over one million users on GitHub, little is known about the issues faced by Ruby developers. This study aims to investigate the key topics, trends, and difficulties faced by Ruby developers by analyzing over $498,000$ Ruby-related questions on Stack Overflow (SO), followed by a survey of $154$ Ruby developers. We employed BERTopic modeling and manual analysis to develop a taxonomy of $35$ topics, grouped into six main categories. Our findings reveal that Web Application Development is the most commonly discussed category, while Ruby Gem Installation and Configuration Issues emerged as the most challenging topic. Analysis of trends on SO showed a steady decline. A survey of $154$ Ruby developers demonstrated that $31.6\%$ of the participants find the Core Ruby Concepts category particularly difficult, while Application Quality and Security is found to be difficult for over $40\%$ of experienced developers. Notably, a comparison between survey responses and SO metrics highlights a misalignment, suggesting that perceived difficulty and objective indicators from SO differ; emphasizing the need for improved metrics to capture developer challenges better. Our study provides insights about the challenges Ruby developers face and strong implications for researchers.
@@ 1. Introduction
This study offers valuable insights into the main challenges that Ruby developers face, paving the path for library contributors and tool developers to create resources that effectively address these pain points and ensure improvements in the Ruby ecosystem that align with developer needs. We open-source our collected dataset and codes to enable the reproducibility of the results. [^1] https://github.com/niktaakbarpour/Unveiling-Ruby
@@ 2. Related Works
Novelty of our work. While the current literature covers developers’ challenges in other programming languages and provides insights into Ruby’s popularity, community engagement, and collaborative practices, it does not identify the specific challenges faced by Ruby developers. Our study fills this gap by analyzing Ruby-related SO questions and conducting a targeted survey of Ruby developers to identify key areas of difficulty and evolving interests.
@@ 3. Methodology
$\alpha$ (Equation 1) measures how relevant is a specific tag in $T$ to Ruby and $\beta$ (Equation 2) measures how significance is a specific tag in $T$. Tags meeting or exceeding defined thresholds for both heuristics were considered significantly relevant to Ruby. We applied thresholds from previous studies (Yang et al., 2023), setting six thresholds for $\alpha$ = {$0.05$, $0.1$, $0.15$, $0.2$, $0.25$, $0.3$} and six for $\beta$ = {$0.005$, $0.01$, $0.015$, $0.02$, $0.025$, $0.03$}. With $7$ years of software development experience, the first author evaluated all $36$ tag configurations by manually inspecting five random questions per tag to assess their relevance to Ruby programming. The best configuration we selected is $\alpha$ = $0.1$ and $\beta$ = $0.01$, aligning with the findings of earlier research (Yang et al., 2023). This refinement yielded the following tag set, $T$ = {ruby, devise, ruby-on-rails-5, rspec, rubygems, sinatra, activerecord, ruby-on-rails-3, nokogiri, rvm, ruby-on-rails, ruby-on-rails-4, hash}.
@@
Step 3: Data preprocessing. We preprocessed our dataset in multiple stages. First, we removed posts with titles and bodies under $50$ characters ($4,398$ posts), as manual inspection showed they lacked meaningful context. For the rest, we concatenated titles and bodies since our manual analysis showed both contained essential keywords. To create embeddings we removed punctuation marks and stop words from NLTK list (Bird et al., 2009). For creating the top $10$ words for each topic, we further removed escape characters, HTML tags, links, code, numbers, and applied lowercasing as used in previous studies (Haque et al., 2020; Yang et al., 2023; Bagherzadeh and Khatchadourian, 2019; Ahmed and Bagherzadeh, 2018; Zhao et al., 2024).
@@
To construct the taxonomy, we used a combination of GPT-4o and a human-in-the-loop approach, a methodology used in prior studies (Shah et al., 2023). The process began by providing GPT-4o with the titles and bodies of $15$ randomly selected questions from each of the $35$ topics, as well as the topic names. GPT-4o was instructed to generate a one-level taxonomy, aiming to create $15$ to $20$ middle categories based on this information and assign names to each of them. Next, two authors manually reviewed the titles and bodies of the $525$ questions (i.e., $15$ per topic), as well as the topic names. After closely examining the taxonomy generated by GPT-4o, the authors refined the groupings, ultimately producing $18$ middle categories. In the final step, we repeated this process to develop the highest level of the taxonomy. GPT-4o was asked to generate $5$ to $10$ higher-level groupings based on $18$ middle categories and $15$ randomly selected questions from each of them (i.e., $270$ posts), which would serve as the main categories. As a result, six main categories were created. To validate the categories, we sampled $525$ posts—exceeding the $384$ needed for $95\%$ confidence with a $5\%$ margin of error. Two authors independently labeled them into six categories, achieving a Cohen’s Kappa of $0.98$, indicating near-perfect agreement. This validation ensured that all Ruby-related questions could fit within our six-category taxonomy.
@@
Step 6: Topic analysis. To assess the popularity and difficulty of topics, we adopted methods from prior research (Yang et al., 2023; Haque et al., 2020). Popularity is evaluated using two metrics: the average number of views for questions within each topic and the average score of questions in each topic. The average number of views shows how many times that question was viewed by people on SO. When a question is useful to the community, users can upvote that question and otherwise, downvote it. The score of each question is calculated by summing the number of upvotes minus the number of downvotes. So, if a question has a high average score it means that the question was very useful to the community. Following previous studies (Yang et al., 2023), our main metric for popularity is the average number of views. The difficulty is assessed through three distinct metrics that were used in previous studies (Yang et al., 2023; Haque et al., 2020). First, we considered the percentage of questions within each topic that do not have an accepted answer. Second, we evaluated the percentage of questions that remain unanswered. Third, we measured the median time required for questions in each topic to receive an accepted answer. Following previous studies (Yang et al., 2023), our main metric for understanding difficulty is the percentage of questions without an accepted answer. The author of the question can mark one of the answers as an accepted answer. Therefore, if a high percentage of questions in a topic do not have accepted an answer, we can understand that the topic is difficult. Additionally, we analyzed the trends related to Ruby questions on SO over the years to understand how they have evolved.
@@
This study was approved by the Behavioral Research Ethics Board of the University of British Columbia.
@@
To validate our survey design, we conducted a pilot study with seven graduate students at the University of British Columbia who had experience with Python but not Ruby. Although they were not the target audience, the goal was to assess the clarity, structure, and usability of the survey to ensure it was well-designed for Ruby developers. The pilot study also helped identify potential issues, refine question wording, and estimate the completion time.
After that, we distributed the full survey via Prolific [^4] https://www.prolific.co, a platform previously used in similar research studies (Reid et al., 2022; Russo, 2022). To ensure data quality, participants were pre-screened for proficiency in Ruby and prior programming experience. Since the survey took approximately $10$ minutes to complete, an attention-check question was included, requiring a correct response for the participant’s submission to be considered valid. The survey was published on Prolific in August 2024, yielding a total of $211$ responses, of which $154$ were approved. Approved participants were automatically paid at a rate of £6 per hour following Prolific’s guidelines. To analyze the survey responses, as (Liang et al., 2024), we report the percentage of participants who selected each option for every topic.
@@ 4.1. RQ1: Topics and Taxonomy
Only $2.59\%$ of the questions were related to Software Architecture and Performance. The questions in this category are related to enhancing the efficiency and design of Ruby applications, designing pattern queries, and improving modularity across multiple Rails applications.
@@ 4.2. RQ2: Analyzing the Topics
Average for all Ruby questions 2092.46 2.49 41.60 9.68 48 Total: 494321 -
@@
To understand whether the more difficult topics are also more popular, we conducted a Kendall rank correlation test at a $95$% confidence level (Hussain and Mahmud, 2019), adopted from prior research (Ahmed and Bagherzadeh, 2018). The null hypothesis ($H_{0}$) is that there is no significant relationship between average popularity and average difficulty, and the alternative hypothesis ($H_{1}$) is that there is a significant relationship between average popularity and average difficulty. The Kendall correlation coefficient was found to be $-0.33$, indicating a moderate negative correlation between popularity and difficulty. However, the $p$-value of $0.47$ suggests that this correlation is not statistically significant. Thus, we cannot confidently assert that a meaningful relationship exists between average popularity and average difficulty.
@@
Figure 3. Trends of six main categories over time. Blue: total number of category-specific questions over the years; Yellow: percentage of category-specific questions relative to all Ruby-related questions.
@@ 4.4. RQ4: Survey Results-Time
On the right side of the chart, we can see the percentage of respondents who reported spending more than $1$ hour to resolve their problems, while the left-hand side reflects the percentages of those who managed to resolve their issues in less than $30$ minutes. The findings indicate that within the Software Architecture and Performance category, $29.5\%$ of participants required more than one hour to address their issues, which is the highest percentage of any category for those taking more than $2$ hours at $13.4\%$. In contrast, the Application Quality and Security category revealed that $44.9\%$ of participants resolved their issues in under $30$ minutes. Furthermore, the Web Application Development category exhibited the highest percentage of participants, at $16.7\%$, who were able to resolve their problems in less than $15$ minutes.
@@
This discrepancy between RQ3 and RQ4 results highlights that developers perceive difficulty and time to resolution differently from what SO data indicates. This could be due to the data from SO and the survey being inherently different and not comparable. Another indication could be similar to the well-known “stated vs. revealed preferences” in behavioral sciences (De Corte et al., 2021); suggesting a potential need for alternative or additional metrics to better capture the complexity and difficulty of issues from a developer’s perspective and SO. As our model shows, the variation in difficulty level is not well explained by time-to-resolution, meaning that time-to-resolution alone may not accurately reflect the difficulty that developers experience. Given the low $R^{2}$ value in our model, this finding likely points to a need for improved or supplementary metrics on SO, as current metrics seem insufficient for capturing perceived difficulty levels.
@@ 5. Discussions
The difference we found in RQ3 and RQ4 results of SO vs. survey data can be attributed to several key factors that highlight the inherent differences in how both data sources measure difficulty. Survey data is subjective, relying on developers’ perceptions that are influenced by personal experiences and prior knowledge among others. In contrast, SO data is based on objective, measurable metrics that capture the broader activity on the platform. As a result, comparing the two becomes challenging, as they measure fundamentally different aspects of the developers’ experiences.